OpenStudy (anonymous):

A silver bar with a mass of 250.0 g is heated from 22.0 º C to 68.5 º C. How much heat does the silver bar absorb?

OpenStudy (anonymous):

q=nc(t2-t1) where n is number of moles and c is heat energy and its value you can find from your book.

OpenStudy (anonymous):

I don't have a book.. I'm homeschooled through online lessons. So could I find it through Google..?

OpenStudy (anonymous):

wait i am giving it to you.

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

here you go http://www2.ucdsb.on.ca/tiss/stretton/database/Specific_Heat_Capacity_Table.html and save that table for later use.

OpenStudy (anonymous):

Okay, I really appreciate that:)

OpenStudy (anonymous):

remember t2=68.5 and t1=22.0

OpenStudy (anonymous):

I think I may have done it wrong.. So this is what I did.. n=6.023*10^23 c=0.240 t2=68.5 t1=22.0 q=(6.023*10^23) * (0.240) * (68.5-22.0) q=(6.023*10^23) * (0.240) * (46.5) q=(1.4455*10^23) * (46.5) q= 6.722 * 10^23 So I'm not sure where I went haywire, but I know it's not right.. Thanks for you help, by the way!

OpenStudy (anonymous):

whats the answer??

OpenStudy (anonymous):

is it 11.16?

OpenStudy (anonymous):

My answer was 6.722*10^23.. but the answers I have to choose from are 0.856kcal, 0.581kcal, 0.275kcal, and 581kcal

OpenStudy (anonymous):

I think where I went wrong was with the moles. I remember something about them always being 6.023*10^23 but not much past that

OpenStudy (anonymous):

do you know about specific heat ?

OpenStudy (anonymous):

I have learned about it, yes

OpenStudy (anonymous):

friend this is also done at here too http://answers.yahoo.com/question/index?qid=20111230111142AAgGboI...... Have a look on that then if you have question then post here.

OpenStudy (anonymous):

It says the question was deleted...

OpenStudy (anonymous):

The specific heat of silver is 0.23 kJ/kg K or 0.057 kcal/kgC The temperature rise = 46.5 C = 46.5 K The mass of silver is 0.25 kg Heat required = 0.23*0.25*46.5 = 2.674 kJ using the conversion 4.18 J = 1 calorie, Heat absorbed = 2.674/4.18 = 0.640 kcal

OpenStudy (jfraser):

your number for "n" is 6.022x10^23, is used improperly. That is the number of things that make up one mole of any substance. The useful equation is\[Q = m*C*\Delta T\] where m is mass, in grams, C is specific heat, and DT is the change in temp. The C of silver is 0.233J/(g*K) or 0.055cal/(g*K). Plug in the appropriate values and compare.

OpenStudy (jfraser):

BTW, no one really uses kcal, all reputable scientific work is done in Joules

OpenStudy (anonymous):

@JFraser Thank you :) & I'm not sure why, but my school does both kcal & Joules both

OpenStudy (anonymous):

I did the equation & I got 639.375 for my answer

OpenStudy (anonymous):

I didn't convert C to K, is that okay? Will the temperature difference be the same?

OpenStudy (jfraser):

yes, the temperature conversion between celsuis and kelvin means that a change of 46.5C is equal to a change of 46.5K. That gets us: \[Q = (250g)*(\frac{0.233J}{g*K})*(46.5K)\]

OpenStudy (jfraser):

I get 2700J, or 2.7kJ

OpenStudy (anonymous):

What would that be in kcal..? & thanks again for helping

OpenStudy (jfraser):

1 calorie is 4.18J, so take 2.7kJ, multiply by 4.18

OpenStudy (anonymous):

11.27 kcal?

OpenStudy (jfraser):

that's what I get. I know that's not one of the answers you have available, but that's what it should be, unless there's some other piece of informatino missing

OpenStudy (anonymous):

Okay, thanks @JFraser & @muhammad9t5 for all your help :) I really appreciate it

OpenStudy (anonymous):

You are welcome always.

OpenStudy (anonymous):

:)